OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone. 8/(9) pi 2pi 8/(3) pi

OpenStudy (blacksteel):

The volume of a cube is side length cubed, so it's 2^3 = 8 The volume of a cone is 1/3 * height * base area = 1/3 * h * pi * r^2 Since the amount of gold in the cube and cone are the same, we know 1/3 * h * pi * r^2 = 1/3 * h * pi * 3^2 = 8 So h * pi * 9 = 24 h*pi = 24/9 = 8/3 h = 8/(3*pi)

OpenStudy (anonymous):

Thanks so much